OpenStudy (anonymous):

Meg conducted a survey of four school cafeterias to find the number of students who like sandwiches for lunch. The results of her survey are recorded in the table below: School Cafeteria Survey School Total Number of Students in the Cafeteria Number of Students Who Like Sandwiches A 38 12 B 48 9 C 26 10 D 24 8 Which school has the greatest percentage of students who like sandwiches for lunch? School A School B School C School D @StudyGurl14

OpenStudy (studygurl14):

Set up these fractions: \(\large\frac{12}{38}\), \(\large\frac{9}{48}\), \(\large\frac{10}{26}\), \(\large\frac{8}{24}\) Turn each one into a decimal to find out which fraction is the largest.

OpenStudy (anonymous):

School C with 38%

OpenStudy (studygurl14):

Correct. :)

OpenStudy (anonymous):

so its c?

OpenStudy (anonymous):

School A has 31% School B has 18% School C has 38% School D has 33%

OpenStudy (anonymous):

Yes

OpenStudy (studygurl14):

yeah

OpenStudy (anonymous):

@ah123456x can u help me with more questions :)

OpenStudy (anonymous):

Ok @geny55

OpenStudy (anonymous):

yay!!!!!!

OpenStudy (anonymous):

What are they?